OpenStudy (anonymous):

What is the product in lowest terms? -8/9 * 5/6

OpenStudy (anonymous):

The same way as before. Show me your work. @TheAsker2002

OpenStudy (anonymous):

Nothing changes even if its negative?

OpenStudy (anonymous):

If it's negative you put a negative sign with it. @TheAsker2002

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

8*5 = 40

OpenStudy (anonymous):

-8*5=-40*

OpenStudy (anonymous):

9*6 = 54

OpenStudy (anonymous):

So you're left with \[\frac{ -40 }{ 54 }\] Right?

OpenStudy (anonymous):

Now simplify it.

OpenStudy (anonymous):

You cant?

OpenStudy (anonymous):

We can. 40 and 54 both are even numbers.

OpenStudy (anonymous):

2 or 4

OpenStudy (anonymous):